Other things remain constant,suppose when the price of good x increase from birr 25 to birr 35 and the quantity demanded of good y increase from 45kg to 90kg.calculate the cross price elasticity of demand between x and y?

2,based on the cross price elasticity above determine whether the two goods are substitute complementary or unrelated goods?

3,interpreted the result


1
Expert's answer
2021-08-29T16:54:24-0400
"E=\frac{90-45}{35-25}\times\frac{35+25}{90+45}=2"

"E>0"

This goods is substituts.